OpenStudy (j.f):

The chance of David winning a game is 1/3. If he plays five times, what is the probability that he will win exactly once?

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

wins once, loses 4 times

OpenStudy (misty1212):

there are 5 different ways he can win once exactly first, second, third, fourth, fifth time so \[5\times \left(\frac{1}{3}\right)\left(\frac{2}{3}\right)^4\]

OpenStudy (j.f):

So this is a Dependent Event

OpenStudy (j.f):

My professor sucks on showing how to properly choose how to set the problem up. The answer is 80/243. Thanks